# Solidification estimates for random walks on supercritical percolation clusters

### v1

@@ 1. Introduction
Figure 1. Schematic illustration of a possible choice of $U_{0}$ with boundary $S=\partial_{\mathscr{S}_{\infty}}U_{0}$ and $\Sigma$ in $\mathscr{S}^{\omega}_{U_{0},\epsilon,\chi}$.
@@
where $\widetilde{\inf}$ is over the same configuration as $\widetilde{\sup}$ in (1.7), and $\mathrm{Cap}^{\omega}(\cdot)$ denotes the random walk capacity on $\mathscr{S}_{\infty}$, see (2.17). On the full lattice $\mathbb{Z}^{d}$, $d\geq 3$ or on $\mathbb{R}^{d}$, $d\geq 3$, uniform lower bounds of the type considered in (1.8) are typically easier to obtain in the context of a convex set $A$ (or its discrete blow-up $A_{N}$), by considering certain a projection attached to $A$ resp. $A_{N}$ (we refer to the discussion in [36, Remark 3.6 (3)] for more on this in the continuum set-up). However, even if $A$ is convex, such a projection is typically ill-defined for $A_{N}\cap\mathscr{S}_{\infty}\subseteq\mathscr{S}_{\infty}$, which reinforces the generality of the methods presented here.
@@ 2.1. Notation
Unless otherwise stated, we assume $d\geq 3$ throughout the article. Occasionally we may also allow $d=2$, especially when we provide examples satisfying the assumptions (see Examples (1), (5), (6)). In fact, part of our approach remains applicable when $d=2$, but substantial modifications to the set-up would be necessary due to the recurrence of the underlying random walk; see also the discussion preceding [36, (2.1)].
@@ 2.2. Model assumptions and examples
    Vacant set of the random walk loop soup in $d\geq 3$. In this case, $\mathbb{P}^{u}=\mathrm{P}^{\overline{\alpha}-u}$, where $\mathrm{P}^{\alpha}$ for $\alpha>0$ is the law of the vacant set $\mathscr{V}^{\alpha}$ of random walk loop soup and $\overline{\alpha}(d)$ is the critical parameter for the strong percolating regime, which can be defined similarly as in Examples (5)(7); see also [2, Theorem 1.3] for its positivity. By Remark 6.6 in [2], the assumptions P1P2, D, and S1S2 are satisfied for $u\in(a,b)$ with $0<a<b<\overline{\alpha}$. It is pausible but open at the moment that $\overline{\alpha}(d)=\alpha_{*}(d)$.
@@ 3. Uniform large-scale behavior over very large regions
$\mathbb{P}^{u}[R_{\mathrm{den}}(\,\cdot\,,\alpha)\geq n]\leq C(\alpha)e^{-c(\alpha)(\log n)^{1+\Delta_{\mathrm{S}}}}.$ (3.4)
In particular, $R_{\mathrm{den}}(\,\cdot\,,\alpha)$ is $\mathbb{P}^{u}$-a.s. finite, and for all $R\geq R_{\mathrm{den}}(\omega,\alpha)$, one has
@@
$\{R_{\mathrm{den}}\geq n\}$ $=\left\{\begin{aligned} &\mbox{there exists $R_{0}\geq n$ and $x\in\mathscr{S}_{\infty}\cap B(0,r_{\alpha,R_{0}})$ such that}\\ &\frac{|\mathscr{S}_{\infty}\cap B(x,R_{0})|}{|B(x,R_{0})|}\in[(1-\alpha)\eta(u),(1+\alpha)\eta(u)]^{c}\end{aligned}\right\}$ (3.6)
@@
$\mathbb{P}^{u}[R_{\mathrm{den}}\geq n]$ $\leq\sum_{R_{0}=n}^{\infty}\sum_{x\in\cap B(0,r_{\alpha,R_{0}})}\mathbb{P}^{u}\left[\frac{|\mathscr{S}_{\infty}\cap B(x,R_{0})|}{|B(x,R_{0})|}\in[(1-\alpha)\eta(u),(1+\alpha)\eta(u)]^{c},x\in\mathscr{S}_{\infty}\right]$
@@
$\mathbb{P}^{u}\left[R_{\mathrm{hk}}(\,\cdot\,,\alpha)\geq n\right]\leq C(\alpha)e^{-c(\alpha)(\log n)^{1+\Delta_{\mathrm{S}}}}.$ (3.9)
In particular, $R_{\mathrm{hk}}(\,\cdot\,,\alpha)$ is $\mathbb{P}^{u}$-a.s.  finite.
@@
Since for $n\geq 1$
$\{R_{\mathrm{hk}}$ $(\,\cdot\,,\alpha)\geq n\}$
@@
$=\bigcup_{R_{0}=n}^{\infty}\bigcup_{x\in B(0,2r_{\alpha,R_{0}})}\{\mathscr{T}(x)\geq R_{0},x\in\mathscr{S}_{\infty}\},$
we can apply a union bound to obtain
@@
    For $\mathbb{P}^{u}$-almost every $\omega$, $\vartheta\in(0,1)$, there exists a finite scale $R_{\mathrm{khk}}(\omega,\alpha,\vartheta)$ and constants $c_{\text{khk}1}(\vartheta)>0$, $0<c_{\text{khk}2}(\vartheta)<1$ such that for all integer $R\geq R_{\mathrm{khk}}(\omega,\alpha,\vartheta)$, the heat kernel $q_{t,B(x_{0},R)\,\cap\,\mathscr{S}_{\infty}}^{\omega}(\cdot,\cdot)$ of the random walk killed upon exiting $B(x_{0},R)\cap\mathscr{S}_{\infty}$ (recall (2.12)) satisfies
@@
[…]
and
@@
We will typically use $\Omega_{\mathrm{reg}}^{\alpha,\vartheta}$ to argue as follows. When the “lowest accessible scale” satisfies $R_{\min}\geq\widetilde{R}$, we obtain uniform estimates over the ball $B(0,r_{\alpha,R_{\min}})$. We choose $\alpha$ to depend on $N$ so that $r_{\alpha_{N},R_{\min}}$ is comparable to or larger than the macroscopic scale. In other words, we can examine the macroscopic sup-norm ball $B(0,r_{\alpha_{N},R_{\min}})$ at any location down to the microscopic scale $R_{\min}$, while retaining good control over the underlying graph structure. Remarkably, the superpolynomial separation between the microscopic and macroscopic scales is sufficient to implement the multi-scale analysis in the proof of the solidification estimates.
Figure 2. Schematic illustration the role of $r_{\alpha,R}$, for $R\geq R_{\min}$. All boxes with center inside $B(0,r_{\alpha,R})\cap\mathscr{S}_{\infty}$ and radius greater or equal to $R$ are well-behaved, but degeneracies (depicted here for the relative volume) occur outside of $B(0,r_{\alpha,R})$.
@@ 4.1. Local density functions
We begin with two properties of the local density functions, which are analogous to the continuum set-up [36, Lemma 1.1] and the discrete set-up with uniformly elliptic conductances [18, Lemma 4.3]. Recall that $\eta(\cdot)$ is assumed to be strictly positive in S2, $\alpha$ is introduced in Proposition 3.1 to represent the deviation between the relative volume of the cluster in a box and its expectation, $R_{\mathrm{den}}(\omega,\alpha)$ is defined in (3.3) as the minimal scale for spatial volume regularity, $\Omega_{\mathrm{den}}^{\alpha}$ is defined in (3.7), and the sequence $(r_{\alpha,N})_{N\geq 1}$ is defined in (3.2) and refers to the radius of a sup-norm ball in which one has uniformly good regularity properties (see Propositions 3.2 and 3.4).
@@
$\leq$ $~\frac{\left||B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|-|B(\pi_{i},2^{\ell})\cap\mathscr{S}_{\infty}|\right|}{|B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|}+\frac{\left||B(\pi_{i},2^{\ell})\cap U_{1}|-|B(\pi_{i},2^{\ell})\cap U_{1}|\right|}{|B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|}.$
@@
We refer to [36, Lemma 1.2] for the proof of the previous lemma. Applying this result with $\mathrm{P}=\mu_{x,\ell}^{\omega}$, $Y=\sigma_{\ell^{\prime}}^{\omega}$, and
@@
We will use the observation to prove the following proposition, which is analogous to [36, Proposition 1.3] and [18, Proposition 4.5], and shows that a random walk has a non-degenerate probability of entering the set where the local density $\sigma_{\ell^{\prime}}^{\omega}$ takes values close to its average over a ball before exiting the latter.
@@
for all $x\in\mathscr{S}_{\infty}\cap B(0,r_{\alpha,2^{\ell^{\prime}}})$.
@@
Then
@@
We now turn to (4.44) and (4.45). First, (4.44) follows by the definition of stopping times (4.41) and a calculation using triangle inequalities similar to (4.51), and we refer to [36, (1.40)] for details. For the lower bound in (4.45),
@@ 4.2. Resonance set
The “canonical” $I^{\prime}$-family as defined in [36, (2.12)] also exists in our case, if we replace the local density function $\widehat{\sigma}_{\ell}$ by $\sigma^{\omega}_{\ell}$, and the conditions $\widehat{\sigma}_{\ell}(X_{S_{i}})=\frac{1}{2}$ by $\sigma_{\ell}^{\omega}(X_{S_{i}})\in[\frac{1}{2}-\frac{1}{2^{\ell_{\min}}},\frac{1}{2}+\frac{1}{2^{\ell_{\min}}}]$. Given a general $I^{\prime}$-family as above, we also define for $1\leq i\leq I^{\prime}$ the stopping times
@@
$\Gamma_{k}(I^{\prime})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sup P^{\omega}_{x}\left[\inf\left\{s\geq S_{0}\,:\,X_{s}\in\operatorname{Res}_{\mathscr{L}_{*},k}\right\}>\max_{1\leq i\leq I^{\prime}}T_{i}\right],$
@@
It is clear by the construction that (4.77a) and (4.77b) holds for $I\geq I_{0}(\epsilon,k+1,J)$. We hence focus on (4.77c) and (4.77d).
@@
$W_{-1}(-e^{-u-1})>-u-1.$ (4.80)
Applying the inequality (4.80) with $u=\log\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}+1$, we obtain
$\left[-\bigg{(}\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}\bigg{)}^{-1}W_{-1}\bigg{(}-\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}\bigg{)}\right]^{2^{k}}\leq\left[\left(\frac{(1+\frac{k-1}{2})2^{k}}{\epsilon/4}\right)\log\frac{(1+\frac{k-1}{2})2^{k}}{\epsilon/4}\right]^{2^{k}}$
$=\left[\left(\frac{(k+1)2^{k+1}}{\epsilon}\right)\log\frac{(k+1)2^{k+1}}{\epsilon}\right]^{2^{k}}\leq I_{0}(\epsilon,k+1).$
@@
$\geq\sum_{x\in\widetilde{U}_{0}}\sum_{x\in\widetilde{U}_{1}}q^{\omega}_{\frac{81}{4}\cdot 4^{\ell},B(x_{0},\frac{9}{2}\cdot 2^{\ell})}(y,x)q^{\omega}_{\frac{81}{4}\cdot 4^{\ell},B(x_{0},\frac{9}{2}\cdot 2^{\ell})}(x,z)\geq C,$
@@
$\leq P_{x_{0}}^{\omega}[H_{\Sigma}>T_{B(x_{0},5\cdot 2^{\ell^{\prime}})}](1-c_{5}(\chi)))^{\mathbbm{1}_{\{\widetilde{\sigma}_{\ell}^{\omega}(x_{0})\in[\widetilde{\alpha},1-\widetilde{\alpha}]\}}}$.
@@
    A particularly important example is supercritical Bernoulli bond percolation on $(\mathbb{Z}^{d},\mathbb{E}^{d})$, $d\geq 3$, where $\mathbb{P}^{u}$ is a product measure of a Bernoulli distribution with success parameter $u$ on the edges, and $u>p_{c}^{\mathrm{bond}}$. A simple inspection of the proof of Theorem 4.1 shows that our results remain valid if the quenched heat kernel bounds (2.13) (together with an integrability condition as in (2.14)) and the large-deviation estimates for the volume (3.1) hold for the corresponding infinite connected component $\mathscr{S}_{\infty}^{\mathrm{bond}}$. The analogues of (2.13) and (2.14) are the main results in [7] (see Theorem 1.1 and (0.5), p. 3026 therein), and (3.1) is found for instance in [37, Theorem 1.2] (or [21, Proposition 11] for an even more quantitative result). We can therefore obtain the equivalent of (4.8) in this situation, namely that for $\chi\in(0,1)$, $(c_{N})_{N\geq 0}$ a sequence of positive real numbers decreasing to 0, and $\mathbb{P}^{u}$-almost every $\omega_{b}\in\{0,1\}^{\mathbb{E}^{d}}$, one has
@@
We first introduce some further notation following [42]. Let $(\ell_{n})_{n\geq 0}$, $(r_{n})_{n\geq 0}$, and $(L_{n})_{n\geq 0}$ be sequences of positive integers, and consider the rescaled lattice $\mathbb{G}_{n}=L_{n}\mathbb{Z}^{d}$ as a graph with an edge between each pair of $\ell^{1}$-nearest neighbors in the latter. For $x\in\mathbb{Z}^{d}$ and $K,s\in\mathbb{Z}_{+}$, we define $Q_{K,s}(x)=x+\mathbb{Z}^{d}\cap[0,KL_{s})^{d}$ and let $\mathscr{C}_{K,s,r}(x)$ be the largest connected component in $\mathscr{S}_{r}\cap Q_{K,s}(x)$ for $r\in\mathbb{Z}_{+}$ (with ties broken arbitrarily), where we recall that $\mathscr{S}_{r}$ is the set of vertices of $\mathscr{S}$ which are in connected components of $\mathscr{S}$ with $\ell^{1}$-diameter greater or equal to $r$. We write $Q_{K,s}$ and $\mathscr{C}_{K,s,r}$ as a shorthand shorthand if $x=0$. We recall below the definition of $(\overline{D},\overline{I},n)$-bad events with respect to two generic sequences of events $\overline{D}=\{\overline{D}_{x,L_{0}}\,:\,L_{0}\geq 1,x\in\mathbb{G}_{0}\}$ and $\overline{I}=\{\overline{I}_{x,L_{0}}\,:\,L_{0}\geq 1,x\in\mathbb{G}_{0}\}$ from [42, Definition 2.1]. For our purposes, it will be sufficient to consider the specific events in (A.1) and (A.2).
@@
Heuristically, if all vertices in in $\mathbb{G}_{s}\cap Q_{K,s}$ are $s$-good, we can iteratively remove “small” cubic regions at each scale, resulting in a subgraph $\mathscr{Q}_{K,s,0}$ of $\mathbb{G}_{0}$ in which all remaining vertices are 0-good, see [42, Fig. 2] for a visualization. Since only small portions of the domain are removed at each step, the connectivity of $\mathscr{Q}_{K,s,0}$ is maintained, and its volume remains comparable to that of $Q_{K,s}$. Hence, we obtain a globally connected component that satisfies the two-sided volume constraints at small scales. This allows us to conclude that the largest connected component in $Q_{K,s}$ meets the announced volume bounds.

### v2

@@ 1. Introduction
Figure 1. Schematic illustration of a possible choice of $U_{0}$ with boundary $S=\partial_{\mathscr{S}_{\infty}}U_{0}$ and $\Sigma$ in $\mathscr{S}^{\omega}_{U_{0},\epsilon,\chi}$. In the picture, $U_{0}$ appears as the intersection between $\mathscr{S}_{\infty}$ and the shaded area, and $S$ consists of the squares at the boundary of the shaded area.
@@
where $\widetilde{\inf}$ is over the same configuration as $\widetilde{\sup}$ in (1.7), and $\mathrm{Cap}^{\omega}(\cdot)$ denotes the random walk capacity on $\mathscr{S}_{\infty}$, see (2.17). On the full lattice $\mathbb{Z}^{d}$, $d\geq 3$ or on $\mathbb{R}^{d}$, $d\geq 3$, uniform lower bounds of the type considered in (1.8) are typically easier to obtain in the context of a convex set $A$ (or its discrete blow-up $A_{N}$), by considering certain a projection attached to $A$ resp. $A_{N}$ (we refer to the discussion in [34, Remark 3.6 (3)] for more on this in the continuum set-up). However, even if $A$ is convex, such a projection is typically ill-defined for $A_{N}\cap\mathscr{S}_{\infty}\subseteq\mathscr{S}_{\infty}$, which reinforces the generality of the methods presented here.
@@ 2.1. Notation
Unless otherwise stated, we assume $d\geq 3$ throughout the article. Occasionally we may also allow $d=2$, especially when we provide examples satisfying the assumptions (see Examples (1), (5), (6)). In fact, part of our approach remains applicable when $d=2$, but substantial modifications to the set-up would be necessary due to the recurrence of the underlying random walk; see also the discussion preceding [34, (2.1)].
@@ 2.2. Model assumptions and examples
    Vacant set of the random walk loop soup in $d\geq 3$. In this case, $\mathbb{P}^{u}=\mathrm{P}^{\overline{\alpha}-u}$, where $\mathrm{P}^{\alpha}$ for $\alpha>0$ is the law of the vacant set $\mathscr{V}^{\alpha}$ of random walk loop soup and $\overline{\alpha}(d)$ is the critical parameter for the strong percolating regime, which can be defined similarly as in Examples (5)(7); see also [2, Theorem 1.3] for its positivity. By Remark 6.6 in [2], the assumptions P1P2, D, and S1S2 are satisfied for $u\in(a,b)$ with $0<a<b<\overline{\alpha}$. It is plausible but open at the moment that $\overline{\alpha}(d)=\alpha_{*}(d)$.
@@ 3. Uniform large-scale behavior over very large regions
$\mathbb{P}^{u}[R_{\mathrm{den}}(\,\cdot\,,\alpha)>n]\leq C(\alpha)e^{-c(\alpha)(\log n)^{1+\Delta_{\mathrm{S}}}}.$ (3.4)
In particular, $R_{\mathrm{den}}(\,\cdot\,,\alpha)$ is a $\mathbb{P}^{u}$-a.s. finite random variable, and for all $R\geq R_{\mathrm{den}}(\omega,\alpha)$, one has
@@
$\{R_{\mathrm{den}}>n\}$ $=\left\{\begin{aligned} &\mbox{there exists $R_{0}\geq n$ and $x\in\mathscr{S}_{\infty}\cap B(0,r_{\alpha,R_{0}})$ such that}\\ &\frac{|\mathscr{S}_{\infty}\cap B(x,R_{0})|}{|B(x,R_{0})|}\in[(1-\alpha)\eta(u),(1+\alpha)\eta(u)]^{c}\end{aligned}\right\}$ (3.6)
@@
$\mathbb{P}^{u}[R_{\mathrm{den}}>n]$ $\leq\sum_{R_{0}=n}^{\infty}\sum_{x\in\cap B(0,r_{\alpha,R_{0}})}\mathbb{P}^{u}\left[\frac{|\mathscr{S}_{\infty}\cap B(x,R_{0})|}{|B(x,R_{0})|}\in[(1-\alpha)\eta(u),(1+\alpha)\eta(u)]^{c},x\in\mathscr{S}_{\infty}\right]$
@@
$\mathbb{P}^{u}\left[R_{\mathrm{hk}}(\,\cdot\,,\alpha)>n\right]\leq C(\alpha)e^{-c(\alpha)(\log n)^{1+\Delta_{\mathrm{S}}}}.$ (3.9)
In particular, $R_{\mathrm{hk}}(\,\cdot\,,\alpha)$ is a $\mathbb{P}^{u}$-a.s. finite random variable.
@@
Recall the definition of the random variables $\mathscr{T}(\cdot)$ above (2.13), and the convention following (2.14). For $n\geq 1$
$\{R_{\mathrm{hk}}$ $(\,\cdot\,,\alpha)>n\}$
@@
$=\bigcup_{R_{0}=n}^{\infty}\bigcup_{x\in B(0,2r_{\alpha,R_{0}})}\{\mathscr{T}(x)\geq R_{0},x\in\mathscr{S}_{\infty}\}.$
Then we can apply a union bound to obtain
@@
    For any $\vartheta\in(0,1)$, there exist constants $c_{\text{khk}1}(\vartheta)>0$, $0<c_{\text{khk}2}(\vartheta)<1$ and a random variable $R_{\mathrm{khk}}(\cdot,\alpha,\vartheta)$ such that for $\mathbb{P}^{u}$-almost every $\omega\in\Omega$, $R_{\mathrm{khk}}(\omega,\alpha,\vartheta)<\infty$, and for all integers $R\geq R_{\mathrm{khk}}(\omega,\alpha,\vartheta)$, the heat kernel $q_{t,B(x_{0},R)\,\cap\,\mathscr{S}_{\infty}}^{\omega}(\cdot,\cdot)$ of the random walk killed upon exiting $B(x_{0},R)\cap\mathscr{S}_{\infty}$ (recall (2.12)) satisfies
@@
$\Omega_{\text{hk}}^{\alpha,\vartheta}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{\omega\in\Omega\,:\,R_{\mathrm{hk}}(\omega,\alpha)<\infty,R_{\mathrm{khk}}(\omega,\alpha,\vartheta)<\infty\}\in\mathscr{F},$ (3.16)
and, with $\Omega_{0}$ and $\Omega_{\mathrm{den}}^{\alpha}$ as in (2.8) and (3.7) respectively,
@@
We will typically use $\Omega_{\mathrm{reg}}^{\alpha,\vartheta}$ to argue as follows. When the “lowest accessible scale” satisfies $R_{\min}\geq\widetilde{R}$, we obtain uniform estimates over the ball $B(0,r_{\alpha,R_{\min}})$. We choose $\alpha$ to depend on $N$ so that $r_{\alpha_{N},R_{\min}}$ is comparable to or larger than the macroscopic scale. In other words, we can examine the macroscopic sup-norm ball $B(0,r_{\alpha_{N},R_{\min}})$ at any location down to the microscopic scale $R_{\min}$, while retaining good control over the underlying graph structure (see Fig. 2 for an illustration). Remarkably, the superpolynomial separation between the microscopic and macroscopic scales is sufficient to implement the multi-scale analysis in the proof of the solidification estimates.
Figure 2. Schematic illustration of the role of $r_{\alpha,R}$, for $R\geq R_{\min}$. All boxes with center inside $B(0,r_{\alpha,R})\cap\mathscr{S}_{\infty}$ and radius greater or equal to $R$ are well-behaved, but degeneracies (depicted here for the relative volume) occur outside of $B(0,r_{\alpha,R})$.
@@ 4.1. Local density functions
We begin with two properties of the local density functions, which are analogous to the continuum set-up [34, Lemma 1.1] and the discrete set-up with uniformly elliptic conductances [18, Lemma 4.3]. Recall that $\eta(\cdot)$ is assumed to be strictly positive in S2, $\alpha$ is introduced in Proposition 3.1 to represent the deviation between the relative volume of the cluster in a box and its expectation, $R_{\mathrm{den}}(\omega,\alpha)$ is defined in (3.3) as the minimal scale for spatial volume regularity, $\Omega_{\mathrm{den}}^{\alpha}$ is defined in (3.7), and the sequence $(r_{\alpha,N})_{N\geq 1}$ is defined in (3.2) and refers to the radius of a sup-norm ball in which one has uniformly good regularity properties (see Propositions 3.2 and 3.4).
@@
$\leq$ $~\frac{\left||B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|-|B(\pi_{i},2^{\ell})\cap\mathscr{S}_{\infty}|\right|}{|B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|}+\frac{\left||B(\pi_{i},2^{\ell})\cap U_{1}|-|B(\pi_{{i+1}},2^{\ell})\cap U_{1}|\right|}{|B(\pi_{i+1},2^{\ell})\cap\mathscr{S}_{\infty}|}.$
@@
We refer to [34, Lemma 1.2] for the proof of the previous lemma. Applying this result with $\mathrm{P}=\mu_{x,\ell}^{\omega}$, $Y=\sigma_{\ell^{\prime}}^{\omega}$, and
@@
We will use the observation to prove the following proposition, which is analogous to [34, Proposition 1.3] and [18, Proposition 4.5], and shows that a random walk has a non-degenerate probability of entering the set where the local density $\sigma_{\ell^{\prime}}^{\omega}$ takes values close to its average over a ball before exiting the latter.
@@
for all $x\in\mathscr{S}_{\infty}\cap B(0,r_{\alpha,2^{\ell^{\prime}}})$, where here and in the following we use the abbreviation $\{\sigma_{\ell^{\prime}}^{\omega}\in[(\sigma_{\ell^{\prime}}^{\omega})_{x,\ell}^{\omega}-\delta,(\sigma_{\ell^{\prime}}^{\omega})_{x,\ell}^{\omega}+\delta]\}$ for $\{y\in\mathbb{Z}^{d}\,:\,\sigma_{\ell^{\prime}}^{\omega}(y)\in[(\sigma_{\ell^{\prime}}^{\omega})_{x,\ell}^{\omega}-\delta,(\sigma_{\ell^{\prime}}^{\omega})_{x,\ell}^{\omega}+\delta]\}$.
@@
We recall the notation related to random walks introduced at the start of Subsection 2.3, up to (2.12). We have
@@
We now turn to (4.44) and (4.45). First, (4.44) follows by the definition of stopping times (4.41) and a calculation using triangle inequalities similar to (4.51), and we refer to [34, (1.40)] for details. For the lower bound in (4.45),
@@ 4.2. Resonance set
The “canonical” $I^{\prime}$-family as defined in [34, (2.12)] also exists in our case, if we replace the local density function $\widehat{\sigma}_{\ell}$ by $\sigma^{\omega}_{\ell}$, and the conditions $\widehat{\sigma}_{\ell}(X_{S_{i}})=\frac{1}{2}$ by $\sigma_{\ell}^{\omega}(X_{S_{i}})\in[\frac{1}{2}-\frac{1}{2^{\ell_{\min}}},\frac{1}{2}+\frac{1}{2^{\ell_{\min}}}]$. Given a general $I^{\prime}$-family as above, we also define for $1\leq i\leq I^{\prime}$ the stopping times
@@
$\Gamma_{k}(I^{\prime})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sup P^{\omega}_{x}\left[\inf\left\{s\geq S_{0}\,:\,X_{s}\in\operatorname{Res}_{(\mathscr{L}_{*},k)}\right\}>\max_{1\leq i\leq I^{\prime}}T_{i}\right],$
@@
As $c_{3}\geq 2$, $\epsilon\in(0,1)$, and $(k+1)2^{k+1}\geq 2$, using the definition (4.74), one has
$\log 2\cdot I_{0}(\epsilon,k+1)^{-1/2^{k}}=\frac{\epsilon}{c_{3}\cdot(k+1)2^{k+1}}\cdot\frac{\log 2}{\log\left(\frac{(k+1)2^{k+1}}{\epsilon}\right)}\leq\frac{\epsilon}{4}.$
Hence (4.77a) holds for $I\geq{I_{0}(\epsilon,k+1)}$. As for (4.77b), we note that $\log(1-c_{2}(J))<0$ and $I^{-1/2^{k}}(\sqrt{I}-1)\geq 1$, since $I\geq I_{0}(\epsilon,2)\geq 256$, using $k\geq 2$. We now focus on (4.77c) and (4.77d).
@@
$W_{-1}(-e^{-u-1})>{2(-u-1),}$ (4.80)
as $u-\sqrt{2u}+1>0$ for $u>0$. Applying the inequality (4.80) with $u=\log\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}+1$, we obtain
$\left[-\bigg(\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}\bigg)^{-1}W_{-1}\bigg(-\frac{\epsilon/4}{(1+\frac{k-1}{2})2^{k}}\bigg)\right]^{2^{k}}\leq\left[{2}\left(\frac{(1+\frac{k-1}{2})2^{k}}{\epsilon/4}\right)\log\frac{(1+\frac{k-1}{2})2^{k}}{\epsilon/4}\right]^{2^{k}}$
$=\left[{2}\left(\frac{(k+1)2^{k+1}}{\epsilon}\right)\log\frac{(k+1)2^{k+1}}{\epsilon}\right]^{2^{k}}\stackrel{{\scriptstyle{c_{3}\geq 2}}}{{\leq}}I_{0}(\epsilon,k+1).$
@@
$\geq\sum_{x\in\widetilde{U}_{0}}\sum_{z\in\widetilde{U}_{1}}q^{\omega}_{\frac{81}{4}\cdot 4^{\ell},B(x_{0},\frac{9}{2}\cdot 2^{\ell})}(y,x)q^{\omega}_{\frac{81}{4}\cdot 4^{\ell},B(x_{0},\frac{9}{2}\cdot 2^{\ell})}(x,z)\geq C,$
@@
$\leq P_{x_{0}}^{\omega}[H_{\Sigma}>T_{B(x_{0},5\cdot 2^{\ell^{\prime}})}]{(1-c_{5}(\chi))}^{\mathbbm{1}_{\{\widetilde{\sigma}_{\ell}^{\omega}(x_{0})\in[\widetilde{\alpha},1-\widetilde{\alpha}]\}}}$.
@@
    A particularly important example is supercritical Bernoulli bond percolation on $(\mathbb{Z}^{d},\mathbb{E}^{d})$, $d\geq 3$, where $\mathbb{P}^{u}$ is a product measure of a Bernoulli distribution with success parameter $u$ on the edges, and $u>p_{c}^{\mathrm{bond}}$. A simple inspection of the proof of Theorem 4.1 shows that our results remain valid if the quenched heat kernel bounds (2.13) (together with an integrability condition as in (2.14)) and the large-deviation estimates for the volume (3.1) hold for the corresponding infinite connected component $\mathscr{S}_{\infty}^{\mathrm{bond}}$. The analogues of (2.13) and (2.14) are the main results in [7] (see Theorem 1 and (0.5), p. 3026 therein), and (3.1) is found for instance in [37, Theorem 1.2] (or [21, Proposition 11] for an even more quantitative result). We can therefore obtain the equivalent of (4.8) in this situation, namely that for $\chi\in(0,1)$, $(c_{N})_{N\geq 0}$ a sequence of positive real numbers decreasing to 0, and $\mathbb{P}^{u}$-almost every $\omega_{b}\in\{0,1\}^{\mathbb{E}^{d}}$, one has
@@
We first introduce some further notation following [42]. Let $(\ell_{n})_{n\geq 0}$, $(r_{n})_{n\geq 0}$, and $(L_{n})_{n\geq 0}$ be sequences of positive integers, and consider the rescaled lattice $\mathbb{G}_{n}=L_{n}\mathbb{Z}^{d}$ as a graph with an edge between each pair of $\ell^{1}$-nearest neighbors in the latter. For $x\in\mathbb{Z}^{d}$ and $K,s\in\mathbb{Z}_{+}$, we define $Q_{K,s}(x)=x+\mathbb{Z}^{d}\cap[0,KL_{s})^{d}$ and let $\mathscr{C}_{K,s,r}(x)$ be the largest connected component in $\mathscr{S}_{r}\cap Q_{K,s}(x)$ for $r\in\mathbb{Z}_{+}$ (with ties broken arbitrarily), where we recall that $\mathscr{S}_{r}$ is the set of vertices of $\mathscr{S}$ which are in connected components of $\mathscr{S}$ with $\ell^{1}$-diameter greater or equal to $r$. We write $Q_{K,s}$ and $\mathscr{C}_{K,s,r}$ as a shorthand if $x=0$. We recall below the definition of $(\overline{D},\overline{I},n)$-bad events with respect to two generic sequences of events $\overline{D}=\{\overline{D}_{x,L_{0}}\,:\,L_{0}\geq 1,x\in\mathbb{G}_{0}\}$ and $\overline{I}=\{\overline{I}_{x,L_{0}}\,:\,L_{0}\geq 1,x\in\mathbb{G}_{0}\}$ from [42, Definition 2.1]. For our purposes, it will be sufficient to consider the specific events in (A.1) and (A.2).
@@
Heuristically, if all vertices in $\mathbb{G}_{s}\cap Q_{K,s}$ are $s$-good, we can iteratively remove “small” cubic regions at each scale, resulting in a subgraph $\mathscr{Q}_{K,s,0}$ of $\mathbb{G}_{0}$ in which all remaining vertices are 0-good, see [42, Fig. 2] for a visualization. Since only small portions of the domain are removed at each step, the connectivity of $\mathscr{Q}_{K,s,0}$ is maintained, and its volume remains comparable to that of $Q_{K,s}$. Hence, we obtain a globally connected component that satisfies the two-sided volume constraints at small scales. This allows us to conclude that the largest connected component in $Q_{K,s}$ meets the announced volume bounds.